A cash register had 5 times as many quarters as nickels, two fewer dimes than nickels, and 30 pennies. All together, the cash register contains $8.50 in change. How many nickles are in the cash register?
Answer by ptaylor(2048)

We'll let x=number of nickels in register
Then 5x=number of quarters
and (x-2)=number of dimes
and 30=number of pennies
Whatever monetary units we deal in (pennies, nickels, dollars, etc.), we have to be consistent. I choose to deal in pennies.
Now we know that the amt of money in nickels (x)(5)plus the amt of money in dimes(x-2)(10)plus the amt of money in quarters (5x)(25) plus the pennies (30) has to equal 850. (Remember, we are dealing in pennies). Thus, our equation to solve is:
5x+10(x-2)+5x(25)+30=850 Simplifying, we get:
5x+10x-20+125x+30=850 further simplifying, we get:
140x=840
x=6 nickels
5x=(5)(6)=30 quarters
x-2=6-2=4 dimes
and, of course, 30 cents
Hope this helps ----ptaylor